A woman is driving her van with speed 50.0 mi/h on a horizontal stretch of road.

(a)

When the road is wet, the coefficient of static friction between the road and the tires is 0.102. Find the minimum stopping distance (in m).

m

(b)

When the road is dry,

μs = 0.598.

Find the minimum stopping distance (in m).

m
Answer #1

Velocity = 50 milh= 22.352 m/s @ acceleration =-119 = 0.9946 -1 m) f² = vetzas Tx S= 22.352² 2X1 s499.61m © az lig=-0.59889.8